Question

The output characteristics of a transistor in common-emitter configuration can be regarded as straight lines connecting the following points

         

   IB =     20 µA        50 µA          80 µA       

 

VCE(V) 1.0 8.0 1.0 8.0 1.0 8.0

 

IC (mA) 1.2 1.4 3.4 4.2 6.1 8.1

 

Plot the characteristics and superimpose the load line for a 1 kΩ load, given that the supply voltage is 9V and the d.c. base bias is 50 μA. The signal input resistance is 800 Ω. When a peak input current of 30 μA varies sinusoidally about a mean bias of 50 μA, determine

(a) The quiescent collector current 

(b) The current gain 

(c) The voltage gain

(d) The power gain


1
Expert's answer
2021-10-07T09:04:12-0400



From Common Emitter configuration, we can develop the load line analysis equation as follows



9 - Ic * 1 - VCE = 0

VCE = 9 - Ic * 1 ................................Equation(1)


We will plot equation(1) with the output characteristics of the transistor as shown in the image attached below.



(a) From graph we see that quiescent Ic = 4.2 mA


(b) Current gain "\beta" = "\dfrac{}{}" "\dfrac{Ic}{I_B}" = "\dfrac{4.2 \ mA}{30\ \ uA}" = 140


(c) Voltage gain AV = "\beta\ *\ \dfrac{Output\ resistance}{Input\ Resistance}" = 140 * "\dfrac{1000}{800}" = 175




(d) Power gain = Av * "\beta" = 140 * 175 = 24500